general solutions of sinx cosx and tanx

Grade:12

3 Answers

Rinkoo Gupta
9 years ago
sinxcosx=1/2.2sinxcox=(sin2x)/2

-1=sin2x=1

-1/2=(sin2x)/2=1/2

-infinity=tanx=infinty

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty

Arun Kumar IIT Delhi
9 years ago
Hi

Arun Kumar, IIT delhi, Btech, Askiitians Faculty
Rinkoo Gupta
9 years ago
sinxcosx=(1/2).2sinxcosx=(sin2x)/2
-1=or<sin2x< or = 1
-1/2 = or < (sin2x)/2<or = 1/2
-infinityor =<tanx< or = infinity
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty